# Quantum Multimodal Contrastive Learning Framework

### v3

@@ Abstract
1 Introduction
@@
In recent years, contrastive learning has emerged as a powerful technique in self-supervised learning, especially for multimodal data. By learning to associate related pairs from different modalities, contrastive learning has shown impressive results in tasks such as image-text alignment and audio-visual correspondence. However, traditional approaches are constrained by classical computing limitations, particularly when dealing with high-dimensional, complex data like EEG. Decoding images from EEG signals presents a promising avenue for studying brain processes, yet it remains difficult due to the low quality and unpredictability of EEG signals. Overcoming these challenges is crucial for advancing our understanding of visual processing and improving EEG-based image decoding and brain-computer interfaces (BCIs).
Earlier attempts to decode images from EEG data were limited by small datasets, which restricted the ability of models to generalize. More recent efforts have utilized larger datasets, often employing methods where images are presented rapidly to elicit brain responses. However, these methods remain constrained by the limitations of classical neural networks. Despite these advancements, current approaches still fall short, underscoring the need for EEG processing techniques that are better suited to the brain’s mechanisms and the unique characteristics of EEG data.
Although quantum methods have not yet outperformed traditional approaches, recent findings in [4] have shown the potential advantages of quantum encoders for EEG processing. Motivated by these findings, we decided to explore the use of quantum techniques in this context. This led us to introduce a Quantum Multimodal Contrastive Learning Framework (QMCL), an early and exploratory attempt to harness the unique properties of quantum computing, such as superposition and entanglement, to improve the learning process across different data modalities. By integrating Variational Quantum Circuits (VQCs) with classical neural networks, our approach can explore a larger solution space, potentially enabling more efficient and effective learning of multimodal representations. While this hybrid quantum-classical framework is still in its early stages, it holds promise for advancing beyond the capabilities of classical methods and opening up new possibilities for applications that were previously out of reach.
Our QMCL framework is an exploratory approach to a relatively challenging task, offering a promising new direction in processing complex data like EEG. This paper contributes the following:
@@
2 Related Works
2.1 Classical Contrastive Learning
Contrastive learning has achieved significant success in unimodal settings. Early models such as SimCLR [5] utilize random data augmentations to generate positive sample pairs and employ cosine similarity for unsupervised learning. Another prominent model, MoCo (Momentum Contrast) [6], introduces a momentum encoder and dynamic dictionary updating, effectively handling large-scale data and outperforming SimCLR in performance. BYOL (Bootstrap Your Own Latent) further advances contrastive learning by removing the need for negative samples, leveraging self-supervised learning to generate high-quality feature representations.
In multimodal contrastive learning, CrossCLR [7] stands out as a significant model that extends contrastive learning to cross-modal representation learning. CrossCLR leverages the correlations between different modalities to improve performance under unsupervised learning. Additionally, CLIP (Contrastive Language-Image Pre-training) [8] marks a milestone in multimodal contrastive learning by simultaneously learning correspondences between text and images, achieving strong multimodal understanding in an unsupervised setting. ALIGN (A Large-scale Image and Noisy Text Embedding) is similar to CLIP, aligning images and text within a shared embedding space using large-scale pre-training, demonstrating the practical potential of multimodal contrastive learning.
These models illustrate the evolution of contrastive learning from unimodal to multimodal applications, highlighting its broad applicability across various data modalities. From unimodal models like SimCLR and MoCo to multimodal models like CLIP and ALIGN, contrastive learning has become a key technique in unsupervised learning, driving advancements in numerous fields. Contrastive learning has shown great success in unimodal settings. Early models like SimCLR use random data augmentations to generate positive pairs and cosine similarity for unsupervised learning. MoCo (Momentum Contrast) introduces a momentum encoder and dynamic dictionary, improving performance on large-scale data. BYOL (Bootstrap Your Own Latent) advances the field by removing the need for negative samples, relying on self-supervised learning for high-quality feature representations.
2.2 Multimodal Contrastive Learning
In recent years, after the success of the traditional contrastive learning models on the same modal data like text and image [9, 10, 11, 12], the development of multimodal contrastive learning has reached significant advancements in the field of self-supervised learning, particularly in tasks that contain the integration of multiple types of data. This method leverages the strengths of various modalities (e.g., text, images, video) to boost model generalization across diverse datasets. Multimodal contrastive learning aligns representations from different modalities within a shared embedding space, facilitating robust, modality-invariant feature learning. This enhances capabilities in cross-modal retrieval and zero-shot learning. Typically, a two-tower network architecture processes each modality independently, with outputs converging in the embedding space where contrastive loss minimizes distances between similar pairs and maximizes distances between dissimilar ones. One of the most popular and successful multimodal contrastive learning framework is CLIP [13], which project both the image and text in to the same feature space. Nevertheless, because datasets containing both time-series signals like EEG and image data are quite rare, there has been little research applying contrastive learning methods to this combination of temporal and visual information. To our best knowledge, [14] is maybe the first work introduced the EEG-image contrastive learning on obtaining the EEG-image representation for image reconstruction downstream task but do not do the zero-shot classification. [15] introduced the EEGClip network for joint representation learning between EEG signal and image but it just do supervised learning. [16] first try to design the EEG encoder on EEG-image contrastive learning , but the work only modified the encoders. This area remains largely uncharted and calls for new, specialized contrastive learning techniques to handle these joint time-series and image modalities effectively.
Decoding visual information from the human brain has long been a challenging task at the intersection of neuroscience and computer science [17, 18, 19, 20]. Although significant progress has been made in understanding static visual inputs, quickly and accurately extracting meaningful information from natural images remains difficult [21, 22]. Early research primarily relied on functional magnetic resonance imaging (fMRI) [23, 24, 25], which has demonstrated the ability to capture meaningful content and structural details from visual processing in the brain. However, fMRI depends on detecting changes in blood oxygen levels, leading to a time lag of several seconds per stimulus, which limits its effectiveness in real-time applications. Moreover, fMRI is expensive and requires large, stationary equipment.
In contrast, electroencephalography (EEG) offers higher temporal resolution, immediate data feedback, and portable, cost-effective hardware, making it a promising candidate for personal brain-computer interface technology. However, current methods using EEG to extract semantic information for image classification have not achieved satisfactory results [26, 27, 16], highlighting the need for improved approaches. Previous methods often relied on supervised learning techniques with a limited set of image categories, ignoring the inherent correlations between visual stimuli and neural responses [27, 28, 15]. These limitations reduce their effectiveness in real-world scenarios that require the generalization to recognize novel, unfamiliar objects. To address these challenges, [29] first attempted zero-shot classification using the largest available EEG-image database [20] with a multilayer MLP and joint EEG-image-text representation, while [16] employed a contrastive learning method. However, [16] used a basic contrastive learning framework based on CLIP [13]. The recent work by [30] improved upon this by using an enhanced neural encoder and an updated CLIP encoder, achieving better results than [16]. Nevertheless, these approaches remain confined to classical machine learning frameworks, and to our knowledge, no similar attempts have been made in the field of quantum machine learning. Our work builds upon the [30] framework and EEG encoder by introducing a self-supervised learning method for EEG-based image decoding and incorporating quantum layers into both the neural and image encoders. This represents a pioneering effort in this field.
2.3 Quantum Machine Learning on Image Data
The integration of quantum computing into machine learning has opened new field for solving complex computational tasks, particularly in the field of image classification. Quantum machine learning (QML) leverages the principles of quantum mechanics, such as superposition and entanglement, to process information in ways that classical systems cannot achieve efficiently. This section explores the application of QML techniques to image data, focusing on the development and evaluation of hybrid quantum neural networks (QNNs) tailored for image classification tasks. [31] introduces two quantum machine learning models designed to address the computational challenges of image classification posed by the growing volume of visual data. One is a hybrid quantum neural network with parallel quantum circuits, outperforming its classical counterpart while using significantly fewer parameters. The other is the model uses a quanvolutional layer to map the input image into lower-resolution feature maps. Despite having fewer parameters, it matches the performance of a classical convolutional neural network. [32] extending the QNN to a scalable variational quantum deep neural network (VQDNN) that supports end-to-end training. That work designing three classifiers for different image recognition tasks using hybrid PCA-VQDNN, amplitude encoding-based VQDNN, and hybrid encoding-based VQDNN to outperform classical neural networks even with a limited number of qubits available in current hardware, suggesting that the approach is promising for practical applications in image recognition using quantum computing. In the survey [33], several recent studies on image classification using quantum machine learning are mentioned. This research direction, compared to traditional machine learning, is still very young but rapidly growing. The integration of quantum computing into machine learning offers new solutions for complex tasks like image classification. Quantum machine learning (QML) utilizes principles like superposition and entanglement to process information more efficiently than classical systems. This section examines the use of QML techniques in image classification, focusing on hybrid quantum neural networks (QNNs). [31] introduces two QML models: a hybrid QNN with parallel quantum circuits that outperforms its classical equivalent with fewer parameters, and a model using a quanvolutional layer to achieve similar performance to classical CNNs with less complexity. [32] extends QNNs to scalable variational quantum deep neural networks (VQDNNs) with end-to-end training, demonstrating superior performance in image recognition tasks even with limited qubits. The survey [33] highlights the rapid growth of this emerging field, despite being in its early stages compared to traditional machine learning.
2.4 Quantum Machine Learning on Time-Series Data
Similar to image data, there have also been recent attempts to apply quantum machine learning to time-series data. The goal is to leverage the properties of quantum information to either reduce the number of model parameters while maintaining the same performance or achieve better data representations. The hybrid quantum-classical VQC algorithm for time-series, in particular, preserves the traditional neural network’s ability to extract temporal and/or spacial structural features while mapping the learned representations to a Hilbert space via quantum layers for feature learning. [34] presents two quantum-classical hybrid machine learning architectures inspired by the multilayer perceptron (MLP). The models were evaluated on both univariate time series (Mackey-Glass time series and USD-to-euro currency exchange rate) and multivariate time series (Lorenz attractor and Box-Jenkins Gas Furnace). The results were competitive with MLP, CNN, and long short-term memory (LSTM) models with a similar number of trainable parameters, suggesting that integrating quantum computing elements could enhance classical models. [35] introduced quantum computing methods for time-series reconstruction and classification, framing these tasks as a quantum unconstrained binary optimization (QUBO) problem solvable by quantum processors. The approach discretizes time-series data, converting it into a set cover problem, enabling efficient reconstruction and semi-supervised classification. The method shows competitive results with classical techniques, using less data, and offers robustness in handling data permutations. [36] introduce a hybrid quantum-classical LSTM model, termed QLSTM. They demonstrate that QLSTM effectively learns various types of temporal data and, in some cases, converges faster or achieves better accuracy than its classical counterpart. The variational approach used in QLSTM reduces the requirements for qubit counts and circuit depth, making it more suitable for implementation on noisy intermediate-scale quantum (NISQ) devices. For the specific EEG application on QML, [4] provided Quantum-EEGNet (QEEGNet), a novel hybrid neural network that integrates quantum computing with the classical EEGNet [37] architecture to enhance EEG encoding and analysis. By incorporating quantum layers, QEEGNet can capture more intricate patterns in EEG data and offer computational advantages.The findings of that work demonstrate the significant potential of quantum-enhanced neural networks in EEG analysis, pointing to promising new directions for research and practical applications.
@@ 2.4 Quantum Machine Learning on Time-Series Data
2.5 Quantum Contrastive Learning
@@
Recent advances in quantum machine learning have focused primarily on unimodal data, exploring quantum-enhanced methods for tasks like image classification and time series prediction. All the works mentioned below concentrate on single-modality data, without addressing the complexities of learning across multiple modalities. [38] introduces a hybrid quantum-classical architecture for self-supervised learning of visual representations. The study reveals a numerical advantage for small-scale QNNs in learning visual representations, and the quantum models perform comparably to classical models on real quantum devices. This paper [39] proposes a model that fuses Supervised Contrastive Learning (SCL) with VQC and employs Principal Component Analysis (PCA) for dimensionality reduction. The approach effectively addresses the challenge of limited training data and demonstrates potential applications in medical image analysis. Experimental results show that the model achieves high accuracy across several medical imaging datasets, especially when using a small number of qubits (e.g., 2 qubits), highlighting the advantages of quantum computing. [40] presents a novel quantum-enhanced supervised contrastive learning model focused on addressing data scarcity issues. Experiments demonstrate that this model achieves significant accuracy in image classification tasks, even with very limited labeled data. The model exhibits stable performance on real quantum devices, proving its adaptability when faced with limited data. This paper [41] proposes a quantum-enhanced self-supervised learning framework for effective mental health monitoring. The framework uses a quantum-enhanced LSTM encoder to improve representation learning for time series data through contrastive learning. The study shows that the model significantly outperforms traditional self-supervised learning models in predictive capability and achieves high F1 scores across multiple datasets.
@@
Table 1: Comparison of the top-1 accuracy results of QSTConv, QSTConv-GA, QNervFormer and QNervFormer-GA on the ThingsEEG test Dataset for 200-way zero-shot classification.
Subject 1 Subject 2 Subject 3 Subject 4 Subject 5 Subject 6
QSTConv $3.2\pm 1.5\%$ $2.1\pm 1.9\%$ $4.5\pm 0.7\%$ $3.5\pm 3.4\%$ $2.0\pm 2.5\%$ $1.8\pm 1.2\%$
QSTConv-GA $4.2\pm 1.4\%$ $3.0\pm 1.0\%$ $3.3\pm 1.3\%$ $6.0\pm 1.2\%$ $2.6\pm 1.4\%$ $3.6\pm 1.1\%$
QNervFormer 4.3±$\pm 1.7\%$ $4.1\pm 1.4\%$ $\textbf{5.6}\pm 1.5\%$ $5.6\pm 1.4\%$ $4.3\pm 0.4\%$ $4.4\pm 0.7\%$
[…]
Subject 7 Subject 8 Subject 9 Subject 10 Average
QSTConv $3.4\pm 2.1\%$ $4.6\pm 1.1\%$ $3.8\pm 1.5\%$ $6.0\pm 0.7\%$ $3.5\pm 1.7\%$
QSTConv-GA $4.9\pm 0.5\%$ $4.4\pm 0.7\%$ $3.9\pm 1.1\%$ $3.9\pm 1.2\%$ $4.0\pm 2.5\%$
QNervFormer $5.5\pm 1.4\%$ $4.8\pm 2.1\%$ $\textbf{5.9}\pm 2.0\%$ $\textbf{6.6}\pm 1.4\%$ $5.1\pm 1.4\%$
QNervFormer-GA $\textbf{6.4}\pm 1.2\%$ $\textbf{5.8}\pm 1.3\%$ $5.6\pm 1.7\%$ $6.1\pm 1.9\%$ $\textbf{5.2}\pm 1.3\%$
@@
Table 2: Comparison of the top-5 accuracy results of QSTConv, QSTConv-GA, QNervFormer and QNervFormer-GA on the ThingsEEG test Dataset for 200-way zero-shot classification.
Subject 1 Subject 2 Subject 3 Subject 4 Subject 5 Subject 6
QSTConv $12.6\pm 4.8\%$ $9.6\pm 6.1\%$ $16.8\pm 1.4\%$ $12.4\pm 11.0\%$ $7.6\pm 6.1\%$ $8.9\pm 5.6\%$
QSTConv-GA $16.7\pm 5.3\%$ $13.4\pm 1.9\%$ $15.0\pm 2.1\%$ $\textbf{24.4}\pm 2.1\%$ $11.4\pm 2.2\%$ $16.7\pm 2.7\%$
QNervFormer $\textbf{18.0}\pm 3.7\%$ $\textbf{18.6}\pm 2.2\%$ $\textbf{21.9}\pm 0.8\%$ $23.1\pm 1.4\%$ $\textbf{15.5}\pm 1.1\%$ $\textbf{20.2}\pm 2.9\%$
QNervFormer-GA $17.4\pm 2.9\%$ $17.7\pm 2.3\%$ $20.2\pm 2.1\%$ $22.4\pm 1.9\%$ $14.5\pm 1.8\%$ $19.5\pm 2.2\%$
Subject 7 Subject 8 Subject 9 Subject 10 Average
QSTConv $13.3\pm 6.0\%$ $20.0\pm 2.4\%$ $16.3\pm 3.3\%$ $19.7\pm 3.9\%$ $13.7\pm 5.0\%$
QSTConv-GA $16.6\pm 2.3\%$ $20.0\pm 2.1\%$ $13.5\pm 1.6\%$ $20.3\pm 2.7\%$ $16.8\pm 2.5\%$
QNervFormer $20.6\pm 2.4\%$ $19.8\pm 8.5\%$ $\textbf{20.3}\pm 5.4\%$ $23.7\pm 1.4\%$ $20.2\pm 3.0\%$
QNervFormer-GA $\textbf{23.0}\pm 2.9\%$ $\textbf{24.5}\pm 2.3\%$ $19.3\pm 3.2\%$ $\textbf{24.8}\pm 2.1\%$ $\textbf{20.3}\pm 2.4\%$
@@
2.6 Image Processing with QCLIP-ViT
@@
Figure 3: The architecture of Quantum-CLIP-Vision transformer (QCLIP-ViT), it uses the pretrained CLIP-ViT weight and project it into qubits’ embedding sopace.
Algorithm 1 Quantum Multimodal Contrastive Learning Framework (QMCL)
1:Input: (Image, EEG) $\triangleright$ stimulus & response
2:Model: $Enc_{img}$: QCLIP-ViT or its variance, $Enc_{eeg}$: QSTConv, QNervFormer or its variance
3:
4:# E : $(batch,channel,electrode,data\ sample)$triangleright batch of input EEGs
5:# I : $(batch,channel,height,width)$triangleright batch of input images
6:
7:# $\tau$ : learned temperature parameter
8:
9:# extract normalized representations from the raw image and EEG
10:$E_{f}$ = Norm(Linear($Enc_{eeg}$(E)))
11:$I_{f}$ = Norm($Enc_{img}$(I)) $\triangleright$ can be obtained before training
12:
13:
14:# scaled pairwise cosine similarity
15:logits = dot($E_{f}$, $I_{f}$.t) $\times$ $e^{\tau}$
16:
17:# symmetric loss function
18:labels = arange(batch) $\triangleright$ self-supervised learning label
19:$loss_{e}$ = CrossEntropyLoss(logits, labels, axis=0)
20:$loss_{i}$ = CrossEntropyLoss(logits, labels, axis=1)
21:$total\_loss$ = ($loss_{e}$ + $loss_{i}$) / 2
2.7 EEG Processing with Quantum Encoders

### v4

@@ Abstract
Index Terms:
electroencephalography (EEG), contrastive learning, quantum machine learning, quantum algorithm, deep learning, brain-computer interface
I Introduction
@@
In recent years, contrastive learning has become a powerful tool in self-supervised learning, especially for multimodal data. By associating related pairs from different modalities, it has achieved impressive results in tasks like image-text alignment and audio-visual correspondence. However, classical approaches struggle with high-dimensional, complex data such as EEG, making image decoding from EEG signals challenging due to the low quality and unpredictability of the signals. Overcoming these challenges is essential for advancing visual processing and improving EEG-based image decoding and brain-computer interfaces (BCIs). Earlier attempts to decode images from EEG data were limited by small datasets, restricting model generalization. While recent efforts with larger datasets and rapid image presentations have made some progress, they are still constrained by classical neural networks. These limitations underscore the need for EEG processing techniques that align better with the brain’s mechanisms and the unique properties of EEG data.
Although quantum methods have yet to surpass traditional approaches, recent findings in [4] suggest that quantum encoders hold promise for EEG processing. Inspired by these results, we developed a Quantum Multimodal Contrastive Learning Framework (QMCL), an exploratory approach that leverages quantum computing’s unique properties, such as superposition and entanglement, to enhance learning across different data modalities. By integrating Variational Quantum Circuits (VQCs) with classical neural networks, our approach explores a broader solution space, potentially enabling more efficient and effective multimodal representation learning. While still in its early stages, this hybrid quantum-classical framework shows promise for surpassing the limitations of classical methods and unlocking new possibilities for complex applications like EEG processing.
Our QMCL framework offers a novel approach to tackling the complexities of EEG data, contributing the following advancements:
@@
II Related Work
II-A Classical Contrastive Learning
In multimodal contrastive learning, CrossCLR [5] extends the approach to cross-modal representation learning by leveraging correlations between different modalities. CLIP [6] marks a major breakthrough, learning text-image correspondences in an unsupervised setting. Similarly, ALIGN [7] aligns images and text within a shared embedding space through large-scale pre-training, showcasing the practical potential of multimodal contrastive learning. These models demonstrate the evolution of contrastive learning from unimodal to multimodal applications, underscoring its versatility across data modalities. From SimCLR [8] and MoCo [9] to CLIP and ALIGN, contrastive learning has emerged as a key technique in unsupervised learning, driving advancements across various fields. [10] has tried to do EEG-image contrastive learning with transformer models, and the work is the state-of-the-art of the task.
II-B Quantum Machine Learning on Time-Series Data
Recent efforts have extended quantum machine learning (QML) to time-series data, aiming to leverage quantum information properties to either reduce model parameters while maintaining performance or improve data representations. Hybrid quantum-classical VQC algorithms for time-series preserve the neural network’s ability to extract temporal and spatial features while mapping learned representations to a Hilbert space via quantum layers. [11] presents two hybrid architectures inspired by multilayer perceptrons (MLP), evaluated on univariate (Mackey-Glass, USD-to-euro) and multivariate (Lorenz attractor, Box-Jenkins Gas Furnace) time series. These models performed competitively with MLP, CNN, and LSTM models of similar complexity, suggesting that integrating quantum elements can enhance classical models. [12] introduced a hybrid quantum-classical LSTM model (QLSTM) that effectively learns temporal data, sometimes converging faster or achieving better accuracy than classical LSTMs. The variational approach in QLSTM reduces qubit and circuit depth requirements, making it suitable for noisy intermediate-scale quantum (NISQ) devices. For EEG applications, [4] proposed Quantum-EEGNet (QEEGNet), a hybrid model integrating quantum computing with the classical EEGNet [13] architecture. QEEGNet enhances EEG encoding by capturing intricate patterns in EEG data, demonstrating the significant potential of quantum-enhanced neural networks in EEG analysis and paving the way for further research and practical applications.
II-C Quantum Contrastive Learning
@@
Recent advances in quantum machine learning have primarily focused on unimodal data, enhancing tasks like image classification and time series prediction. All the following works concentrate on single-modality data without addressing the complexities of multimodal learning. [14] introduces a hybrid quantum-classical architecture for self-supervised visual representation learning, demonstrating the numerical advantages of small-scale QNNs and their comparable performance to classical models on real quantum devices. [15] proposes a model that combines Supervised Contrastive Learning (SCL) with VQC and uses Principal Component Analysis (PCA) for dimensionality reduction, achieving high accuracy in medical image analysis, especially with a small number of qubits. [16] presents a quantum-enhanced supervised contrastive learning model addressing data scarcity, showing significant accuracy in image classification tasks with limited labeled data and stable performance on real quantum devices. [17] proposes a quantum-enhanced self-supervised learning framework for mental health monitoring, utilizing a quantum-enhanced LSTM encoder that significantly outperforms traditional models in predictive capability, achieving high F1 scores across multiple datasets. Despite these advancements, all existing works are confined to unimodal data. In contrast, our work introduces Quantum Multimodal Contrastive Learning (QMCL), the first approach to leverage quantum computing for learning across multiple data modalities. This breakthrough allows for richer and more complex representations, opening new avenues for applications that require the integration of diverse data sources.
@@
TABLE I: Comparison of the top-1 accuracy results of QSTConv, QSTConv-GA, QNervFormer and QNervFormer-GA on the ThingsEEG test Dataset for 200-way zero-shot classification.
Subject 1 Subject 2 Subject 3 Subject 4 Subject 5 Subject 6 Subject 7 Subject 8 Subject 9 Subject 10 Average
QSTConv 3.2% 2.1% 4.5% 3.5% 2.0% 1.8% 3.4% 4.6% 3.8% 6.0% 3.5%
QSTConv-GA 4.2% 3.0% 3.3% 6.0% 2.6% 3.6% 4.9% 4.4% 3.9% 3.9% 4.0%
QNervFormer 4.3% 4.1% 5.6% 5.6% 4.3% 4.4% 5.5% 4.8% 5.9% 6.6% 5.1%
QNervFormer-GA 3.9% 4.2% 5.4% 6.1% 4.4% 4.5% 6.4% 5.8% 5.6% 6.1% 5.2%
@@
TABLE II: Comparison of the top-5 accuracy results of QSTConv, QSTConv-GA, QNervFormer and QNervFormer-GA on the ThingsEEG test Dataset for 200-way zero-shot classification.
Subject 1 Subject 2 Subject 3 Subject 4 Subject 5 Subject 6 Subject 7 Subject 8 Subject 9 Subject 10 Average
QSTConv 12.6% 9.6% 16.8% 12.4% 7.6% 8.9% 13.3% 20.0% 16.3% 19.7% 13.7%
QSTConv-GA 16.7% 13.4% 15.0% 24.4% 11.4% 16.7% 16.6% 20.0% 13.5% 20.3% 16.8%
QNervFormer 18.0% 18.6% 21.9% 23.1% 15.5% 20.2% 20.6% 19.8% 20.3% 23.7% 20.2%
QNervFormer-GA 17.4% 17.7% 20.2% 22.4% 14.5% 19.5% 23.0% 24.5% 19.3% 24.8% 20.3%
III Methodology
@@
III-A Image Processing with QCLIP-ViT
@@
III-B EEG Processing with Quantum Encoders